OpenStudy (iesmee):

Solve for the z coordinate of the system x+2y+5z=-1 2x-y+z=2 3x+4y-4z=14

OpenStudy (texaschic101):

x + 2y + 5z = -1 -->(-2) 2x - y + z = 2 ------------------ -2x - 4y - 10z = 2 (result of multiplying by -2) 2x - y + z = 2 ------------------add -5y - 9z = 4 2x - y + z = 2 ---->(3) 3x + 4y - 4z = 14 -->(-2) ------------------ 6x - 3y + 3z = 6 (result of multiplying by 3) -6x - 8y + 8z = -28 (result of multiplying by -2) -----------------add -11y + 11z = - 22 -5y - 9z = 4 --->(-11) -11y + 11z = -22 -->(5) -------------------- 55y + 99z = -44 (result of multiplying by -11) -55y + 55z = - 110 (result of multiplying by 5) ------------------add 154z = - 154 z = -1 -5y - 9z = 4 -5y - 9(-1) = 4 -5y + 9 = 4 -5y = 4 - 9 -5y = -5 y = 1 x + 2y + 5z = -1 x + 2(1) + 5(-1) = -1 x + 2 - 5 = -1 x - 3 = -1 x = -1 + 3 x = 2 check... 2x - y + z = 2 2(2) - 1 - 1 = 2 4 - 2 = 2 2 = 2 (correct) x = 2, y = 1, and z = -1